Elaan paddled out to check on her crab traps. Going to the traps she caught a falling tide, that increased her normal speed by 2mph. but coming back it decreased her speed by 2 mph. Going with the tide , the trip took only 10 min. The return, against the tide, took 20 min. How far out are the traps from the shore?
I've tried to set up the chart and declare the variables but I've only seen current problems where the the unknown is the speed of the current/boat. I don't know how to set up the system to find the distance. Help Please =) Thank you for your time.

Answer by mananth(16949)

let boat speed = x
with current = x+2 mph ----- time = 10 minutes
against current = x-2 mph----time = 20 minutes
Distance = same both ways
r*t = distance
1/6(x+2)= 1/3(x-2)
3(x+2)=6(x-2)
3x+6=6x-12
6x-3x=18
3x=18
x=6
D= rt
D= 1/6 *8= 4/3 miles
D=rt
D= 1/3*4= 4/3 miles
Distance was 4/3 miles